OpenStudy (anonymous):

Find g '(x) for g(x) = sec(2x). a) g '(x) = 2sec^2(2x) b) g '(x) = 2sec(2x) c) g '(x) = 2sec(2x)tan(2x) d) g '(x) = 2tan^2(2x) I think i got: C? i just want someone to check my answer :)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i think its c

OpenStudy (jhannybean):

This involves using the chain rule: \(\frac{d}{dx}(\sec(x) \cdot \frac{d}{dx}(x)\)

OpenStudy (anonymous):

2sec(2x)tan(2x)

OpenStudy (anonymous):

c is the answer

OpenStudy (anonymous):

that's what i got too @laetitiacarmel :) and @Jhannybean ok, do u think its correct as C? :)

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Do you know your derivatives @mathrulezz ?

OpenStudy (anonymous):

Yeah, chain rule is the way to go, lets look at it, g(x) = sec(2x) g'(x) = sec(2x)tan(2x)*(2x)'

OpenStudy (anonymous):

is someone taking calculus 2?

OpenStudy (anonymous):

Derivative of secx = secxtanx

OpenStudy (jhannybean):

And yes, C looks correct.

OpenStudy (anonymous):

ok thank u guys :)

OpenStudy (anonymous):

I also used prime notation, even though the one Jhanny showed you is the "legit" way, but this is easier to see what's going on especially if you're just starting derivatives.